5. The Friendly National Bank holds $50 million in reserves at its Federal Reserve District Bank. The required reserves ratio is 12 percent.
a. If the bank has $600 million in deposits, what amount of vault cash would be needed for the bank to be in compliance with the required reserves ratio?

b. If the bank holds $10 million in vault cash, determine the required reserves ratio that would be needed for the bank to avoid a reserves deficit.
c. If the Friendly National Bank experiences a required reserves deficit, what actions can it take to be in compliance with the existing required reserves ratio?

Answer by jim_thompson5910(35256)

a)
12% of 600 million = 0.12*600 = 72 million

So they need $72 in cash in the vault to be in compliance with the required reserves ratio

---------------------------------------------------------------------

b)

(10/600)*100 = 0.01666666666667*100 = 1.666666666667 = 1 and 2/3% reserves ratio

So the answer is 1.6667% or 1 and 2/3%

--------------------------------------------------------------------
c)

It needs to add more cash to the vault or lower the total amount of deposits in the bank.

If it goes with option 1, then it needs to have $72 million (so it needs to add 72 - 10 = 62 million)


If it goes with option 2, then it should reduce the total amount deposited to 10/0.12 = 83.333 million (roughly). So it would need to remove 600 - 83.333 = 516.667 million from its deposits (and give that money back to the customers)


Either option is viable and it will depend on the situation which one is best.